# A mathematical theory for understanding when abstract representations emerge in neural networks

### v1

@@ IV Abstract representation emerges in the hidden layer, independent of single-neuron nonlinearity
Our central strategy is to (i) prove these results for a perturbed version of the nonlinear function, and then (ii) extend the result via a continuity argument. We outline these key arguments for the single‐element class case ($n=1$) with whitened input. Detailed derivations, as well as the scenarios with multi‐element classes ($n\geq 2$) and target‐aligned inputs, are provided in SI §3.
@@ IV.1 Threshold nonlinear activation
where $E_{r}(\cdot;\rho)$ is given by Eq.(III.1) except that the regularization strength $\lambda_{1}$ is rescaled $\lambda_{1}\rightarrow\lambda_{1}B_{0}^{-2}$. This result is proved via the method of majorization and Schur-convexity [9, 58] (SI §3.1-3.2).
Note that the argument $\phi^{\delta}(\mathbf{h})$ in $E_{r}(\phi^{\delta}(\mathbf{h});\rho)$ takes values from a hypercube $\phi^{\delta}(\mathbf{h})\in[0,\phi_{+}(+\infty)]^{P}$ within the nonnegative orthant $\mathbb{R}^{P}_{\geq 0}$. Minimizing $E_{r}(\phi^{\delta}(\mathbf{h});\rho)$ (SI §3.2), we find that the optimal representation kernel is still of the form $K^{\delta}[\rho_{*}]=b_{*}(d_{Y}\mathbf{1}\mathbf{1}^{T}+K_{Y})$, with $b_{*}$
@@
Since the optimal kernel (and $b_{*}$) is independent of $\delta>0$, we can carry out the limiting argument $\delta\downarrow 0^{+}$ and find that this optimal kernel $K^{\delta}[\rho_{*}]=b_{*}(d_{Y}\mathbf{1}\mathbf{1}^{T}+K_{Y})$ indeed remains unchanged even if $\delta=0$ (SI §3.4).
@@ IV.2 Odd-symmetric nonlinear activation
Specifically, the unique optimal kernel is found to be
@@
$\lambda_{1,2}$ are small enough to ensure $b_{*}\geq 0$. We note that there is no constant shift in the kernel.
Moreover, the set of optimal preactivations (SI §3.3) can now vary continuously within a subspace, having the form
@@
Despite the differences, the optimal kernel in Eq. (35) still represents an abstract representation. Finally, repeating the limiting argument as before shows that this optimal representation kernel remains unchanged when taking $\delta\downarrow 0^{+}$ (SI §3.4).
@@ IV.3 Differences in single-neuron tuning
On the other hand, for odd activation functions [Eq. (33)] (including the linear activation function), the optimal preactivations [Eq. (36)] can freely rotate in a subspace (when $\mathbf{h}$ is small). So in this case, neurons in the hidden layer generally exhibit mixed selectivity [78, 29] (Fig. 7F).
@@ V Extensions
We first derive some general properties of the optimal representation kernel derived using our analytical framework (Section II). Given any input and output matrices $X,Y$, denote the set of optimal representation kernels as $S(X,Y)\subseteq\mathbb{R}^{P\times P}$. According to Eq. (II.3), the optimal preactivation matrix $H$ and thus the optimal representation kernel only depend on the input and output kernels, $S(X,Y)\equiv S(K_{X},K_{Y})$. Moreover, this set is invariant under global scaling factors of the input and output kernels (for ReLU), $S(\alpha_{X}K_{X},\alpha_{Y}K_{Y})=S(K_{X},K_{Y})$ if $\alpha_{X},\alpha_{Y}>0$.
(1) Denote the set of optimal measures of Eq. (15) as $\mathcal{P}(K_{X},K_{Y})\subseteq M_{+}(X)$. Since the problem is a convex optimization problem, the optimal solution set, $\mathcal{P}(K_{X},K_{Y})$ is a convex set and in particular, is simply connected. Because the kernel $K[\cdot]$ is a linear map on $M_{+}(X)$, the set of optimal representation kernel matrices $S(K_{X},K_{Y})=K[\mathcal{P}(K_{X},K_{Y})]$ is also convex and simply connected. For ReLU nonlinearity, $S(K_{X},K_{Y})$ is a subset of the set of completely positive matrices and can be attained when the network width $M\geq 1+P(P+1)/2$ [85]. This property is usually called mode connectivity of the loss landscape [28, 4, 64, 86], and our analytical framework provides an alternative way to uncover this property.
(2) For input kernel $K_{X}$ and any optimal kernel $K[\rho_{*}]\in S(K_{X},K_{Y})$, denote the set of optimal preactivations as $A(K_{X},K_{Y})$. Now if a rank one perturbation is added to the input kernel, $K_{X}^{new}=K_{X}-a\mathbf{v}\mathbf{v}^{T},\,a\in\mathbb{R},\mathbf{v}\in\text{Range}K_{X}$, and moreover,
@@
(3) If the pseudo-inverse of the input kernel $K_{X}^{\dagger}$ has negative off-diagonals, we can show that the optimal set of neural preactivations lies entirely within the nonnegative orthorant, $A(K_{X},K_{Y})\subseteq\mathbb{R}^{P}_{\geq 0}$. And this property holds independently of the output kernel $K_{Y}$. Under this condition, the problem of training two-layer network [Eq. (II.1)] is equivalent to a dictionary learning problem with positive representations. Input kernels $K_{X}$ satisfying such properties are known as inverse M-matrix [27, 38, 39, 19]. (A notable special case are strictly ultrametric matrices [62, 26], see details in SI §4.2).
(4) For any optimal measure $\rho_{*}\in\mathcal{P}(K_{X},K_{Y})$, the output of the network for a new data point $\mathbf{x}\in\mathbb{R}^{d_{X}}$ is (SI §4.3),
@@
where $X$ is the augmented input matrix [Eq. (5)] and $Y$ is the output matrix [Eq. (II.1)]. This result thus allows us to investigate the generalization property of the optimal network solution. As an illustration, we apply it to simple compositional‐generalization tasks (see SI §4.4).
@@ V.1 Anisotropic input-output geometry
According to the definition of $\mathbf{v}_{i}$ (Section II.1), this would correspond to anisotropic output labels where the $i$th output direction is scaled by a factor $d_{i}>0$ for different $i$’s, yielding a hierarchical structure in the output kernel (Fig. 8A). We assume that the inputs are also anisotropic ($c_{i}>0$’s are different) and are target-aligned,
@@
This recovers scenarios in previous sections if all $c_{i}$’s and $d_{i}$’s are equal to each other.
We solve the mean-field problem [Eq. (II.4)] in this case (SI §4.5). And this yields the unique optimal hidden representation kernel
@@
When $d_{i}=1,c_{i}=c_{Y}$, this solution recovers Eq. (30). We see that up to a global translation, the effect of anisotropic input and output just scales the $i$th direction of the hidden representation by a factor $b_{i}^{*}$ that depends on $d_{i}$. The hidden representation still has a hypercube (or hyper-rectangular) geometry (Fig. 8C) and is abstract. However, the anisotropy in the training data induces a more pronounced stage-like transition in the learning dynamics (Fig. 8C).
@@ V.2 Deep neural network
We are interested in when the last layer in this network exhibits abstract representations (Fig.8D) given small regularization parameters $\lambda_{l}\ll 1,l=1,2,..,L$.
@@
Finally, the analytical framework is extended to analyze the optimal representation in recurrent neural networks (SI §4.7). The loss functional in this case can be written as a functional over the space of measures of the temporal trajectory of the neural preactivations. As in the deep feedforward network case, this loss functional is generally not a convex functional. The KKT condition for a recurrent network depends on the full temporal trajectory of preactivations (SI §4.7), rather than factorizing cleanly at each layer as in the feedforward case (SI §4.6). Nevertheless, we find that for whitened inputs and the data model in Section II.1, the learned representation at the last timestep remains abstract (Fig. 8E).
@@ VI Discussion
The low-dimensional disentangled abstract representations of task-relevant variables have been observed in multiple brain areas of multiple species (Fig.1, [8, 91, 24, 70, 61]).Despite the ubiquity of this type of representational geometry, we still do not know the network mechanisms that lead to its formation. Here, we showed that abstract representations naturally emerge in simple feedforward neural networks when the networks are optimized on tasks that depend on the variables that should be represented in an abstract format. In these networks, the geometry of the hidden layer reflects the geometry of the outputs (labels, Fig.1-2). The input geometry or input encoding also has an important effect on the representational geometry learned in the hidden layer (Fig.4C, Fig.6D-F and [2, 41]): when the input and output geometries are aligned and abstract, it is not surprising that the hidden representation will also be abstract. It is less obvious that whitened inputs — despite not being aligned with any specific low-dimensional representation geometry — nonetheless facilitate the emergence of low-dimensional abstract neural representations. This facilitation arises from the maximal dimensionality of the whitened data, which allows the hidden-layer neural representation to ”move around more freely” in the high-dimensional space and inherit the low-dimensional structure from the output. In the brain, these high-dimensional representations are likely the result of some form of ’recoding’ [57, 56] which is probably implemented in the hippocampus[30, 7, 75, 13, 90]. Earlier work has shown that even random dimensionality expansion increases the separability between different stimuli and helps discrimination [59, 5, 52]. Here, we suggest that these expansion layers offer another benefit: they facilitate better representation learning in downstream brain areas. These learned low-dimensional representations would support sample-efficient generalization for novel tasks [6, 93].
@@
In order to investigate the optimal neural representation in the network models, we developed a mathematical framework that maps the original problem of weight space optimization to an optimization problem over the neural preactivations [Fig.3A-C, Eq. (II.3)]. The new problem corresponds to an effective model where neurons interact with each other through a ”mean-field” induced by the overall distribution of neural activity [Eq. (14), Fig. 3]. We derive the KKT condition for the optimal distribution of neural activity [Eq. (II.4)]. For a large class of input kernels, the KKT condition is equivalent to a nonnegative quadratic programming problem, which we solve in closed form when the outputs exhibit cubic symmetry (SI §2.2). Crucially, because the new optimization problem is convex, any solution of the KKT condition is automatically a global minimum of the loss function. We show that, in many cases, all such solutions correspond to an abstract neural representation. To our knowledge, this provides the first theoretical result showing the robust emergence of abstract representation in nonlinear neural networks, when trained to binary decision tasks closely related to prior neuroscience experiments (Section II.1, [8, 17]). These results, together with the analytical framework developed to derive them, constitute the main innovations of this work.
@@
Our work thus provides insights into why abstract representations appear in the brain [8, 91, 24, 70, 61]. Our analytical framework is also a more general tool for analyzing the optimal representations for various tasks used in the machine learning literature. Our framework is a complementary approach to many existing works on analyzing learning in two-layer neural networks. Here we highlight and compare it with two sets of related work (see a more comprehensive discussion in SI §5):
@@
A substantial body of prior work has investigated two-layer networks from the perspective of Bayesian neural networks. Leveraging tools from statistical physics, these studies have shown that certain scaling limits of the model lead to Gaussian equivalence properties [31, 49, 71], wherein the neural preactivations follow a joint Gaussian distribution—a regime closely linked to the kernel limit or lazy regime [97, 23, 36]. More recent work has explored alternative scaling limits [98, 99, 10, 94, 48] that are connected to the mean-field limit of deep networks. While these results primarily concern the (infinite-width or infinite-input-dimension) scaling limits, our framework enables direct analysis of optimal solutions in finite-width networks with finite-dimensional inputs. These results for finite-width networks allow us to make statements about the optimal representation in various infinite-width limits (SI §5.1). Furthermore, whereas Bayesian approaches typically average the network properties over the posterior distribution, our method provides insights into the structure of individual global minima of the loss function. We comment that when deriving the scaling limit from our finite-width results, we first take the infinite temperature limit $\beta\rightarrow+\infty$ followed by infinite-width limit $M\rightarrow+\infty$. As a consequence, the resulting network always learns the low-dimensional features in the data. This is different from the usual Bayesian network setting where $M\rightarrow+\infty$ is taken first and then $\beta\rightarrow+\infty$, and feature learning only happens when using a special weight scaling in the loss function.
@@
Another series of work examined the learning dynamics of two-layer networks. Earlier works focused on single-layer perceptron [84, 22, 73] and deep linear networks [82, 83]. Recently, the learning dynamics of two-layer nonlinear networks were also studied both in the mean-field regime [60, 15, 79, 87] and in finite-width settings [11]. These analyses on nonlinear networks are typically only tractable for a one-dimensional output. However, abstract representations (Fig. 1) concern the relationship between the hidden representations of different output dimensions (see Section II.2). So we adopt a different approach here: rather than tracking the entire training trajectory, we directly analyze global minima of the loss function [Eq. (II.1)]. An intriguing future direction would be to extend the existing results for learning dynamics with one-dimensional outputs, to those tasks in Section II.1 that require multi-dimensional outputs, and investigate how the hidden representation kernel evolves during training.
@@
A byproduct of our analysis is a characterization of single-neuron selectivity in the optimal network solution. This is determined by the set of optimal preactivation vectors $A(K_{X},K_{Y})$ [Eq.(23), (25) and (36)], whose components specify the neural response to each stimulus in the training set. A longstanding question in neuroscience is whether neurons exhibit “interpretable” tuning—activity explained by a single task-relevant variable—or “mixed” tuning, where responses depend on combinations of multiple variables [29]. Experimental data and models have suggested that the tuning type depends on details of task structure [41, 96, 21], brain regions [74], and species. For the task and network architecture studied here (Section II.1), we find that the nonlinearity in the hidden layer plays a key role (see also [2]): depending on its form, the task optimization yields either distinct neural modules tuned to individual binary latent variables or neurons with linear mixed selectivity to multiple latent variables. In the brain, the neuron nonlinearity varies across regions due to differences in single-neuron biophysical and morphological properties [76].
@@
Despite such variability on the single-neuron level, we demonstrated that in the large-network limit, the emergence of abstract representations on the population level is robust to the specific form of single-neuron nonlinearities. This universality result offers a potential explanation for recent experimental observations of abstract representations across diverse brain areas. Previous studies have shown that neural network models, under certain scaling limits, exhibit Gaussian universality or Gaussian equivalence properties, typically attributed to central limit theorem-like mechanisms [33, 31]. In contrast, the optimal preactivation distributions ($\rho_{*}$) in our model are non-Gaussian, suggesting a different origin of universality. We believe that the shared task structure [Section II.1] on which all these models are trained is the underlying reason for this universal abstract representation. This insight aligns with recent empirical findings that networks with different architectures, when trained on similar tasks, often converge to similar neural representations—a phenomenon referred to as the Platonic representation hypothesis [50, 45, 34]. Our results thus provide a tractable mathematical model for this hypothesis and illustrate that neural representations for a task on the population level (in terms of representation kernel), can be relatively insensitive to single-neuron-level response details.
@@
Finally, the order parameter $\rho$—defined as the distribution over neural preactivations—provides a powerful tool for exploiting the permutation symmetry inherent in these network models [4, 60, 87, 15]. We anticipate that this formulation will not only aid in analyzing the optimal solutions in the networks studied here but may also be extended to investigate a broader class of permutation-symmetric models, including ResNet and transformer models in modern AI systems, as well as to analyze learning dynamics in network models trained with biological learning rules.

### v2

@@ IV Abstract representation emerges in the hidden layer, independent of single-neuron nonlinearity
Our central strategy is to (i) prove these results for a perturbed version of the nonlinear function, and then (ii) extend the result via a continuity argument. We outline these key arguments below for the single‐element class case ($n=1$) with whitened input. Detailed derivations, as well as the scenarios with multi‐element classes ($n\geq 2$) and target‐aligned inputs, are provided in SI §3.
@@ IV.1 Threshold nonlinear activation
where $E_{r}(\cdot;\rho)$ is given by Eq. (III.1) except that the regularization strength $\lambda_{1}$ is rescaled $\lambda_{1}\rightarrow\lambda_{1}B_{0}^{-2}$. This result can be proved via the method of majorization [8, 57] (SI §3.1-3.2).
Note that the argument $\phi^{\delta}(\mathbf{h})$ in $E_{r}(\phi^{\delta}(\mathbf{h});\rho)$ takes values from a hypercube $\phi^{\delta}(\mathbf{h})\in[0,\phi_{+}(+\infty)]^{P}$ within the nonnegative orthant $\mathbb{R}^{P}_{\geq 0}$. Minimizing $E_{r}(\phi^{\delta}(\mathbf{h});\rho)$ is a nonconvex copositive programming problem similar to the ReLU case (SI §3.2), and we find that the optimal representation kernel is still of the form $K^{\delta}[\rho_{*}]=b_{*}(d_{Y}\mathbf{1}\mathbf{1}^{T}+K_{Y})$, with $b_{*}$
@@
Note that the optimal kernel (and $b_{*}$) is independent of $\delta>0$, so we can carry out the limiting argument $\delta\downarrow 0^{+}$ and find that the same kernel $K^{\delta}[\rho_{*}]=b_{*}(d_{Y}\mathbf{1}\mathbf{1}^{T}+K_{Y})$ is still optimal even if $\delta=0$ (SI §3.4).
@@ IV.2 Odd-symmetric nonlinear activation
Specifically, the unique optimal kernel in this case is found to be
@@
$\lambda_{1,2}$ are small enough to ensure $b_{*}\geq 0$. We note that there is no constant shift (the $\mathbf{1}\mathbf{1}^{T}$) in the above kernel.
Moreover, the set of optimal preactivations (SI §3.3) forms a subspace, having the form
@@
Despite the differences, the optimal kernel in Eq. (35) still represents an abstract representation. Finally, repeating the limiting argument as before shows that this representation kernel remains optimal when taking $\delta\downarrow 0^{+}$ (SI §3.4).
@@ IV.3 Differences in single-neuron tuning
On the other hand, for odd activation functions [Eq. (33)] (including the linear activation function), the optimal preactivations [Eq. (36)] can freely rotate within a subspace (when $\mathbf{h}$ is small). So in this case, neurons in the hidden layer generally exhibit mixed selectivity [74, 27] (Fig. 7F), where many components of $\mathbf{h}$ are nonzero.
@@ V Extensions
We first discuss some general properties of the optimal representation kernel derived using our analytical framework (Section II). Given any input and output matrices $X,Y$, denote the set of optimal representation kernels as $S(X,Y)\subseteq\mathbb{R}^{P\times P}$. According to Eq. (II.3), the optimal preactivation matrix $H$ and thus the optimal representation kernel only depend on the input and output kernels, $S(X,Y)\equiv S(K_{X},K_{Y})$. Moreover, this set is invariant under global scaling factors of the input and output kernels (for $\phi=$ ReLU and small $\lambda_{1,2}$), $S(\alpha_{X}K_{X},\alpha_{Y}K_{Y})=S(K_{X},K_{Y})$ if $\alpha_{X},\alpha_{Y}>0$.
(1) Denote the set of optimal measures of Eq. (15) as $\mathcal{P}(K_{X},K_{Y})\subseteq M_{+}(X)$. Since the problem is a convex optimization problem, the optimal solution set, $\mathcal{P}(K_{X},K_{Y})$ is a convex set and in particular, is simply connected. Because the kernel $K[\cdot]$ is a linear map on $M_{+}(X)$, the set of optimal representation kernel matrices $S(K_{X},K_{Y})=K[\mathcal{P}(K_{X},K_{Y})]$ is also convex and simply connected. For ReLU nonlinearity, $S(K_{X},K_{Y})$ is a subset of the set of completely positive matrices and by Caratheodory’s theorem, can be attained when the network width $M\geq 1+P(P+1)/2$ [84]. This property is usually called mode connectivity of the loss landscape [26, 3, 63, 85], and our analytical framework provides an alternative way to uncover this property.
(2) For input kernel $K_{X}$, denote $K[\rho_{*}]\in S(K_{X},K_{Y})$ as the optimal kernel with the set of optimal preactivations $A(K_{X},K_{Y})$. Now if a rank one perturbation is added to the input kernel, $K_{X}^{new}=K_{X}-a\mathbf{v}\mathbf{v}^{T},\,a\in\mathbb{R},\mathbf{v}\in\text{Range}K_{X}$, and moreover, its magnitude and direction satisfy
@@
(3) If the pseudo-inverse of the input kernel $K_{X}^{\dagger}$ has negative off-diagonals, we can show that the optimal set of neural preactivations lies entirely within the nonnegative orthorant, $A(K_{X},K_{Y})\subseteq\mathbb{R}^{P}_{\geq 0}$. And this property holds independently of the output kernel $K_{Y}$. In this specific scenario, the problem of training two-layer network [Eq. (II.1)] is equivalent to a dictionary learning problem with positive representations. Input kernels $K_{X}$ satisfying such properties are known as inverse M-matrix [25, 38, 37, 17] (SI §4.2).
(4) For any optimal measure $\rho_{*}\in\mathcal{P}(K_{X},K_{Y})$, the output of the network for a new data point $\mathbf{x}=(\mathbf{x}_{data},1)^{T}\in\mathbb{R}^{d_{X}+1}$ is (SI §4.3),
@@
where $X$ is the augmented input matrix [Eq. (5)] and $Y$ is the output matrix [Eq. (II.1)]. This result thus allows us to investigate the generalization property of the optimal network solution. As an illustration, we apply it to a simple compositional‐generalization task (see SI §4.4).
@@ V.1 Anisotropic input-output geometry
According to the definition of $\mathbf{v}_{i}$ (Section II.1), the above output kernel would correspond to anisotropic output labels where the $i$th output direction is scaled by a factor $d_{i}>0$ for different $i$’s, yielding a hierarchical structure (Fig. 8A). We assume that the inputs are also anisotropic ($c_{i}>0$’s are different) and are target-aligned,
@@
This kernel recovers scenarios in previous sections if $d_{i}=1,c_{i}=c_{Y}$ for all $i$’s.
We solve the mean-field problem [Eq. (II.4)] in this case and find that the unique optimal hidden representation kernel is (SI §4.5)
@@
We see that up to a global translation, the effect of anisotropic input and output simply scales the $i$th direction of the hidden representation by a factor $b_{i}^{*}$ that depends on $d_{i}$. The hidden representation still has a hypercube (or hyper-rectangular) geometry (Fig. 8C) and is abstract. However, the anisotropy in the training data induces a more pronounced stage-like transition in the learning dynamics (Fig. 8B).
@@ V.2 Deep neural network
We are interested in when the last layer of the network exhibits abstract representations (Fig. 8D) for small regularization parameters $\lambda_{l}\ll 1,l=1,2,..,L$.
@@
Finally, the analytical framework is extended to analyze the optimal representation in recurrent neural networks (SI §4.7). The loss in this case can be written as a functional over the space of measures of the temporal trajectory of the neural preactivations. As in the deep feedforward network case, this loss functional is generally not convex. The KKT condition for a recurrent network depends on the full temporal trajectory of preactivations (SI §4.7), rather than factorizing cleanly at each layer as in the feedforward case (SI §4.6). Nevertheless, we find that for whitened inputs and the data model in Section II.1, the learned representation at the last timestep remains abstract (Fig. 8E).
@@ VI Discussion
The low-dimensional disentangled abstract representations of task-relevant variables have been observed in multiple brain areas of multiple species (Fig.1, [7, 90, 23, 65, 61]). Despite the ubiquity of this type of representational geometry, we still do not know the network mechanisms that lead to its formation. Here, we showed that abstract representations of the latent variables naturally emerge in simple feedforward neural networks when the networks are optimized on tasks whose output (i.e. behavior readout) depend on those latent variables. In these networks, the geometry of the hidden layer reflects the geometry of the outputs (labels, Fig. 1-2). The input geometry or input encoding also has an important effect on the representational geometry learned in the hidden layer (Fig. 4C, Fig. 6D-F and [1, 40]): when the input geometry aligns with the abstract output geometry, it is not surprising that the hidden representation will also be abstract (Section III.3). It is less obvious that whitened inputs (Section III.1-III.2), despite not being aligned with any specific low-dimensional representation geometry, nonetheless facilitate the emergence of low-dimensional abstract neural representations. One possible intuition is that this facilitation arises from the maximal dimensionality of the whitened data, which allows the hidden-layer neural representation to ”move around more freely” in the high-dimensional space and inherit the low-dimensional structure from the output. In the brain, these high-dimensional representations are likely the result of some form of ’recoding’ [56, 55], which is probably implemented in the hippocampus [28, 6, 71, 12, 89]. Earlier work has shown that dimensionality expansion increases the separability between different stimuli and helps discrimination [59, 4, 51]. Here, our result suggests that these expansion layers offer another benefit: they facilitate better representation learning in downstream brain areas. These learned low-dimensional representations would support sample-efficient generalization for novel tasks [5, 93].
@@
In order to investigate the optimal neural representation formed in the network, we developed a mathematical framework that maps the original problem of weight space optimization to an optimization problem over the neural preactivations [Fig. 3A-C, Eq. (II.3)]. The new problem corresponds to an effective model where neurons interact with each other through a ”mean-field” induced by the overall distribution of neural activity [Eq. (14), Fig. 3]. We derive the KKT condition for the optimal distribution of neural activity [Eq. (II.4)]. For a large class of input kernels, the KKT condition is equivalent to a nonnegative quadratic programming problem, which we solve in closed form when the output geometry is abstract (SI §2.2). Crucially, because the new optimization problem is convex, any solution of the KKT condition is automatically a global minimum of the loss function. We show that, in many cases, all such solutions correspond to an abstract neural representation. To our knowledge, this provides the first theoretical result showing the robust emergence of abstract representation in nonlinear neural networks, when trained to binary decision tasks closely related to prior neuroscience experiments (Section II.1, [7, 15]). These results, together with the analytical framework developed to derive them, constitute the main innovations of this work.
@@
Our work provides insights into why abstract representations appear in the brain [7, 90, 23, 65, 61]. More broadly, our analytical framework provides a general tool for characterizing optimal representations across a wide range of tasks studied in the machine learning literature. It offers a complementary perspective to existing approaches for analyzing learning in two-layer neural networks. Below, we compare our approach with two major streams of related work (see a more comprehensive discussion in SI §5):
@@
A substantial body of prior work has investigated two-layer networks from the perspective of Bayesian neural networks. Leveraging tools from statistical physics, these studies have shown that certain scaling limits of the model lead to Gaussian equivalence properties [29, 48, 67], wherein the neural preactivations follow a joint Gaussian distribution, a regime closely linked to the kernel or lazy regime [96, 22, 35]. More recent work has explored alternative scaling limits [98, 97, 94, 47] that are connected to the mean-field limit of deep networks. While these results primarily concern the (infinite-width or infinite-input-dimension) scaling limits, our framework enables direct analysis of optimal solutions in finite-width networks with finite-dimensional inputs. The results for finite-width networks allow us to derive results on the optimal representation in various infinite-width limits (SI §5.1). Furthermore, whereas Bayesian approaches typically average the network properties over the posterior distribution, our method provides insights into the structure of individual global minima of the loss function. We comment that when deriving the scaling limit from our finite-width results, we first take the infinite temperature limit $\beta\rightarrow+\infty$ followed by infinite-width limit $M\rightarrow+\infty$. As a consequence, the resulting network always learns the low-dimensional features in the data. This is different from the usual Bayesian network setting where $M\rightarrow+\infty$ is taken first and then $\beta\rightarrow+\infty$, where the feature learning only happens when using a special weight scaling in the loss function.
@@
Another series of work examined the learning dynamics of these networks. Earlier works focused on perceptron [83, 21, 69], deep linear networks [80, 81] and narrow two-layer network in the teacher-student setup [77, 79, 78]. Recently, the learning dynamics of overparametrized two-layer nonlinear networks were studied both in the mean-field regime [60, 14, 75, 86, 9] and in finite-width settings [10]. These recent analyses on nonlinear networks are typically only tractable for a one-dimensional output. However, abstract representations (Fig. 1) concern the relationship between the hidden representations of different output dimensions (Section II.2). So we adopt a different approach here: rather than tracking the entire training trajectory, we directly analyze global minima of the loss function [Eq. (II.1)]. An intriguing future direction would be to extend the existing results for learning dynamics with one-dimensional outputs, to those tasks in Section II.1 that require multi-dimensional outputs, and investigate how the hidden representation kernel evolves during training.
@@
A byproduct of our analysis is a characterization of single-neuron selectivity in the optimal network solution. This is determined by the set of optimal preactivation vectors $A(K_{X},K_{Y})$ [Eq.(23), (25) and (36)], whose components specify the neural response to each stimulus in the training set. A longstanding question in neuroscience is whether neurons exhibit “interpretable” tuning—activity explained by a single task-relevant variable, or “mixed” tuning, where responses depend on combinations of multiple variables [27]. Experimental data and models have suggested that the tuning type depends on details of task structure [40, 95, 19, 20], brain regions [70], and species. For the task and network architecture studied here (Section II.1), we find that the nonlinearity in the hidden layer plays a key role (see also [1]): depending on its form, the task optimization yields either distinct neural modules tuned to individual binary latent variables or neurons with mixed selectivity to multiple latent variables. In the brain, the neuron nonlinearity varies across regions due to differences in single-neuron biophysical and morphological properties [72], which may explain the heterogeneous single-neuron selectivity observed in experimental data.
@@
Despite such variability on the single-neuron selectivity, we demonstrated that in the large-network limit, the emergence of abstract representations on the population level is robust to the specific form of single-neuron nonlinearities. This universality result offers a potential explanation for recent experimental observations of abstract representations across diverse brain areas. Previous studies have shown that neural network models, under certain scaling limits, exhibit Gaussian universality or Gaussian equivalence properties, typically attributed to central limit theorem-like mechanisms [32, 29]. In contrast, the optimal preactivation distributions ($\rho_{*}$) in our model are non-Gaussian, suggesting a different origin of universality. We believe that the shared task structure [Section II.1] on which all these models are trained is the underlying reason for this universal abstract representation. This insight aligns with recent empirical findings that networks with different architectures, when trained on similar or related tasks, often converge to similar neural representations, a phenomenon referred to as the Platonic representation hypothesis [49, 44, 33]. Our results thus provide a tractable mathematical model for investigating this hypothesis and show that neural representations for a task on the population level (in terms of representation kernel), can be relatively insensitive to single-neuron-level response details.
@@
Finally, the order parameter $\rho$, defined as the distribution over neural preactivations, provides a powerful tool for exploiting the permutation symmetry inherent in these network models [3, 60, 86, 14]. We anticipate that this formulation will not only aid in analyzing the optimal solutions in the networks studied here but may also be extended to investigate a broader class of permutation-symmetric problems, including ResNet and transformer models in modern AI systems, as well as to analyze learning dynamics in network models trained with biological learning rules.